Grade 12Algebra

There are 102 teachers in a school of 3300 students. Find the ratio of the number of teachers to the number of students.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given
Number of teachers in a school = 102
Number of students in a school = 3300
Ratio of number of teachers to the number of students = 102 / 3300
= (2 × 3 × 17) / (2 × 3 × 550)
= 17 / 550

Thanks